### v2

# Learning to Predict Program Execution by Modeling Dynamic Dependency on Code Graphs

Abstract
Predicting program behavior without execution is an essential and challenging task in software engineering. Traditional models often struggle to capture dynamic dependencies and interactions within code. This paper introduces a novel machine learning-based framework called CodeFlow, which predicts code coverage and detects runtime errors through Dynamic Dependencies Learning. Utilizing control flow graphs (CFGs), CodeFlow represents all possible execution paths and the relationships between different statements, offering a comprehensive understanding of program behavior. It constructs CFGs to depict execution paths and learns vector representations for CFG nodes, capturing static control-flow dependencies. Additionally, it learns dynamic dependencies through execution traces, which reflect the impacts among statements during execution. This approach enables accurate prediction of code coverage and identification of runtime errors. Empirical evaluations show significant improvements in code coverage prediction accuracy and effective localization of runtime errors, surpassing current models.
I Introduction
Large language models (LLMs) excel in understanding source code and descriptive texts [1, 2, 3, 4, 5, 6, 7]. Their ability to recognize patterns, syntax, and semantics makes them effective at tasks such as code completion, bug detection, and generating human-readable explanations. However, state-of-the-art LLMs [8, 9] exhibit weaknesses in predicting dynamic program behavior, such as code coverage prediction and run-time error detection, etc, which typically require a program executable, but ideally, we want a model to predict them correctly without execution. This limitation arises from their reliance on static code representations, which fail to capture dynamic program behavior and state changes at runtime. Consequently, the models’ token-based predictions result in a superficial understanding of code, lacking context for variable states and control flow across multiple iterations. As a result, they struggle to accurately simulate loops, conditional branches, and the cumulative effects between statements. This shortcoming is further exacerbated by their inability to understand dynamic dependencies and interactions between various statements, making them ill-equipped to handle intricate control flows.
To address these limitations, several approaches have been proposed. The pre-trained model TRACED [10] relies exclusively on the final execution of the last line within a loop to finalize the program states via variable value ranges, which leads to inadequate handling of condition and iteration statements. TRACED employs a variable coverage learning approach, labeling variable occurrences within an executed line. This may fail to capture branching behavior in scenarios where a branch lacks variable occurrences (e.g., having statements like return, exit, etc.) or in cases where a variable occurrence in a true branch occurs in one iteration but not in another. In contrast, CodeExecutor [9], uses UniXcoder [11] on pre-training data including the source code, input values, and the full execution trace with values at each execution step. It heavily relies on UniXcoder to transform the source code and its input into the entire sequence for the execution trace.
Toward dynamic program behavior prediction, we introduce CodeFlow, a code coverage prediction model designed to predict code coverage given source code and its input. CodeFlow leverages a control flow graph (CFG) that helps it better understand and predict the dynamic dependencies in code, including the execution of different branches and loops under varying conditions. With CFGs, we model loops as circular paths, allowing messages to pass through all possible paths and return to the loop node. This captures the aggregate effect of all iterations, ensuring the model understands cumulative changes in variables. To enhance coverage prediction, we focus on learning dynamic dependencies via execution paths on CFG with respect to input values. The CFG provides a detailed representation of the execution flows, capturing intricate paths through sequential, branching, and iterations.
To show CodeFlow’s usefulness, we leverage it to build a tool to statically detect runtime errors in (in)complete code snippets. Platforms like Stack Overflow (S/O) are invaluable resources for developers facing technical issues. However, S/O code snippets may contain hidden defects, runtime errors, and security vulnerabilities, posing potential risks to applications that integrate them. Integrating a code snippet into a codebase for analysis is time-consuming and risky. To mitigate these risks, it is essential to directly analyze online code snippets to reason about their behaviors. However, the key challenge is the incompleteness nature of online code snippets. To do so, we use an LLM to act as a fuzzer, generating inputs to detect runtime errors in a given code snippet. Each input is used in the code snippet, which is then fed into CodeFlow to predict the code coverage. If the code coverage stops unexpectedly and never reaches an exit point, CodeFlow will locate the error.
We conducted an empirical evaluation on CodeFlow. Our findings indicate that it significantly improves code coverage prediction, runtime error detection, and bug localization compared to existing models. Specifically, CodeFlow achieves an accuracy of 75.24% in matching code coverage exactly, outperforming GPT-4o at 68.13%. For branch coverage, CodeFlow reaches 87.88%, significantly higher than GPT-4o’s 78.75%. In runtime error detection, CodeFlow attains a high accuracy of 97.51%, far exceeding the performance of other models. Additionally, CodeFlow maintains high accuracy even on incomplete code snippets, demonstrating its generalization capability. Furthermore, CodeFlow proves highly effective in supporting fuzz testing, particularly in scenarios involving incomplete code snippets where traditional execution is not feasible.
In brief, this paper makes the following contributions:
  • CodeFlow: Dynamic Dependencies Learning for Code Coverage Prediction: A novel code coverage prediction model leveraging CFGs to capture both static and dynamic code dependencies. CodeFlow models loop as circular paths and learning dynamic dependencies among statements.
  • Effective Runtime Error Detection and Localization: CodeFlow analyzes code coverage continuity within CFGs to accurately detect and localize runtime errors.
  • Comprehensive Empirical Evaluation: Experiments show that CodeFlow outperforms existing models in code coverage prediction, runtime error detection, and localization.
II Motivation
II-A Example and Observations
TrueCodeExecutorLLMCode Snippet1>>>>>>>>>x = 102>>>>>>>>>while x > 4:3>>>>>>>>> if x % 2 == 0:4>>>>>>>>> print(’x is even’)5>>>>>>>>> x -= 26$>$ else:7$>$ print("x is odd")8$>$ x -= 19>>>>>>for i in range(100):10>>>>>> x += i11>>>>>>if x % 3 == 0:12 print("x devide by 3 is 0")13$>$elif x % 3 == 1:14$>$ print("x devide by 3 is 1")15$>$else:16$>$ print("x devide by 3 is 2")
Figure 1: Code Coverage Prediction Comparison
Let us use an example to explain the problem and motivate our proposed solution. Fig. 1 shows an example in Python with the input value $x$ = 10 and the code coverage, where ”$>$” indicates the lines of code that are executed during actual execution. We employed two state-of-the-art approaches in CodeExecutor [9] and GPT-4 [8] to predict code coverage for our example and the results are shown in Fig. 1.
II-A1 Observation 1. Conditional Statements
Conditional statements (if-elif-else) present a challenge for existing approaches. For instance, after a For loop, the program checks the value of $x$. LLMs may skip necessary checks, resulting in incorrect predictions, such as jumping directly to the else statement and bypassing elif. This occurs due to their lack of state tracking across lines of code and inability to understand the dependencies between nested branches.
II-A2 Observation 2. Complex Loop Branching
Loops, such as while, contain multiple branches determined by intermediate values of variables, leading to various outcomes such as skipping, entering, exiting, or continuing the loop. These values can change during the loop’s iterations, making accurate prediction difficult with a top-to-bottom approach. For example, in the code snippet (Fig. 1), the while loop on line 2 processes even and odd values of x differently. GPT-4 struggles to grasp the nuances of loop execution due to their reliance on token-based predictions without understanding the dynamic dependencies among statements via the state changes.
Conversely, CodeExecutor [9] correctly skips the else branch by understanding the intermediate values of x throughout the execution trace. However, it performs poorly with complex code requiring multiple iterations to update variable values. Error propagation frequently occurs, leading to incorrect code coverage predictions.
II-A3 Observation 3: Information Loss in Repeated Loops
For loops that run numerous iterations, predicting the outcome based on line-by-line variable states often leads to incorrect results. For example, the for loop on line 11 runs 100 iterations, updating x each time (lines 12-13). While the LLM correctly predicts that the loop will execute all iterations, it fails to understand the cumulative effect on x. This is because the LLM only processes static information, whereas the for loop requires information from the last line in the body to be fed back into the loop. After the loop, the value of x is used to decide which branch of the if-elif-else statement (lines 14-19) will be executed. The LLM incorrectly predicts the outcome because it does not account for the aggregated change in x. Consequently, it fails to accurately simulate the dynamic changes in variable states over multiple iterations.
Figure 2: Control Flow Graph for code in Fig. 1
II-A4 Observation 4: Runtime Error Detection
Detecting runtime errors requires understanding both the static and dynamic dependencies within the code. To determine whether a line contains a runtime error, it is crucial to know which lines are related and affect it (static dependencies). Additionally, understanding how variable changes impact the execution flow (dynamic dependencies) is essential. Existing models struggle with this task because they often fail to capture these intricate dependencies. They do not adequately analyze how changes in variable states influence subsequent lines of code, leading to missed detections. This lack of comprehensive dependency analysis makes it challenging for these models to pinpoint the exact line causing the runtime error and understand its context.
II-B Key Ideas
From the above observations, we design our solution CodeFlow with the following design strategies:
II-B1 Key Idea 1. [Learning Code Execution on Control Flow Graph]
Instead of reasoning the predicted execution on source code, we leverage a graph-based representation for such prediction and code coverage prediction: Control Flow Graph.
Definition 1 (Control Flow Graph - CFG).
A control flow graph (CFG) is a graphical representation of the control flow within a program. Nodes in the graph represent basic blocks of code, such as individual or groups of statements that are executed sequentially, while edges represent the flow of control between these blocks, typically based on conditions such as loops, conditional statements (e.g., if-else), or function calls.
Fig. 2 displays the corresponding CFG of the code in Fig. 1. The CFG illustrates the sequence of execution of statements or code blocks within a program and the conditions that decide the control flow between different blocks, which are divided according to the program semantics. Learning execution on CFG provides several benefits. First, CFGs (Fig. 2) explicitly represent the sequential nature of condition checks, ensuring all paths are considered and the model can accurately predict the execution flow based on all possible conditions. Second, training a model on code coverage using a CFG offers significant advantages over training on source code alone. This allows the model to better understand and predict the dynamic behavior of code, including how different branches and loops are executed based on varying conditions. In contrast, source code only provides static information without context on how the execution evolves. Third, using CFGs to model loops as circular paths, allowing messages to pass through all possible paths and return to the loop node. This method captures the aggregate effect of all iterations, ensuring the model comprehensively captures the cumulative changes in variables.
II-B2 Key Idea 2. [Dynamic Dependencies Learning via Execution Paths on CFG]
A CFG is like a map that provides a blueprint for all possible paths, while an execution path is like a specific travel route on that map, tailored to a particular input of the program. To better predict code coverage, we aim to learn dynamic dependencies among statements on CFG via a large number of execution paths with respect to different inputs. That allows a model to better learn the representations of the execution flows, capturing the dynamic dependencies through sequential, branching, and iteration statements.
II-B3 Key Idea 3. [Detecting Runtime Error via CFG]
Once we have the code coverage, we combine it with the static dependency information between each line of the CFG to detect runtime errors. By checking the continuity in code coverage on the CFG, we can effectively identify a runtime error where the execution path unexpectedly terminates. Specifically, if the predicted path does not reach the EXIT node, we trace back to the furthest node reached without an outgoing edge on the CFG. This node is likely to be the crash point and hence contains the error. By ensuring that the model predicts a continuous path in the CFG, we can accurately detect runtime errors and precisely localize the line containing the bug.
III Approach Overview
Figure 3: CodeFlow: Predictive Code Coverage and Runtime Error Detection with Dynamic Dependencies Learning on CFG
Putting together our above ideas, we develop CodeFlow, a code coverage prediction model that is given a source code and its input and predicts the corresponding code coverage.
Fig. 3 illustrates CodeFlow’s overall architecture. The input is the source code that needs to be predicted.
Step 1. CFG Building (Section IV)
First, the given source code is parsed to build the Abstract Syntax Tree (AST) and the CFG. The input of the source code is encoded as the assignments of the input variables with their values. Additionally, we apply processing steps to generalize and standardize the CFG including normalization of node labels, removal of redundant nodes, and simplification of complex structures.
Step 2. Source Code Representation Learning (Section 4)
The goal of this module is to learn the vector representations (embeddings) for the nodes in the CFG that takes into account the static control-flow dependencies between the statements represented by the connected nodes. We used a Gated Recurrent Unit (GRU) Networks to transform the code into the embeddings that preserve the contextual and semantic information. The output of this step is the CFG structure with each node represented by its corresponding embedding.
Step 3. Dynamic Dependencies Learning (Section VI)
The goal of this step is to learn the dynamic dependencies that pertain to the execution of two connected nodes/statements. Let us call them execution-based dynamic dependencies or dynamic dependencies for short, which indicates whether a statement (represented by a node) would be executed if a node/statement connected with that node is executed. To teach our model on such dynamic dependencies, we use the actual execution traces for the source code and inputs in the training data. We leverage a specialized message-passing scheme with a binary soft decision branching technique to effectively learn the interactions and dependencies that influence code coverage. The output of this step is the CFG structure with its nodes represented by the new vector representations that capture dynamic dependencies during execution.
Step 4. Code Coverage Prediction via Classification (Section VII-A)
The goal of this step is to predict the code coverage for the statements or branches in the given code. Specifically, we use the learned embeddings from Dynamic Dependencies Learning to classify whether a specific node or branch will be covered during actual code execution.
Step 5. Runtime Error Detection and Localization (Section VII-B)
Finally, we use the code coverage predictions from Step 4 to detect whether the code contains runtime errors or not. Moreover, by analyzing the predicted code coverage along with the CFG, we identify nodes where the execution unexpectedly terminates, indicating a potential runtime error.
IV Control Flow Graph Building
In the initial step, we create a CFG from a given code snippet to capture the static dependencies between different code blocks. Fig. 2 illustrates the CFG of the code shown in Fig. 1. However, the original CFG often contains redundant information and lacks clarity in certain nodes, such as those representing loop conditions. To ensure consistency in loop representation and make the CFG easier to process and learn on, we convert for loops into while loops, treating them as condition nodes. This transformation helps maintain only two types of nodes in the graph: operation nodes, which have only one outgoing edge, representing a sequential order of statements, and condition nodes, which have more outgoing edges, representing branching based on conditions, simplifying the embedding process. Additionally, we enhance CFG’s clarity by removing redundant information from each node, such as condition symbols (if, elif, else, and while), and adding markers T to distinguish True and False branches of condition statements. These modifications ensure each node has a uniform structure, making it easier for our model to learn and capture comprehensive information from the code.
In the end of the process, given initial program $P$ will be broken down into graph $G=(V,E)$ with set of node $V=\{n_{1},n_{2},\cdots,n_{L}\}$ with $L\geq 1$ as the number total of nodes. And the edge set $E$ consists of two edge types: forward edge and backward edge. In addition, each node $n_{i}$ also consists of a code statement to represent the semantic information.
V Source Code Representation Learning
Figure 4: Source Code Representation Learning
The fundamental part of modeling the dynamic execution of a program is to statistically analyze the interactions of statements. Capturing the semantic information of those statements is the first stage of almost all machine learning (ML) approaches toward treating code as a sequence of tokens. Following this direction, we treat the node’s statement $n_{i}$ as sequence of lexically tokens $c_{i_{1}},c_{i_{2}},\cdots,c_{i_{L}}$. Each token is then embedded into a vector $x_{i_{t}}=W_{e}c_{i_{t}}$ using a randomly initialized embedding matrix $W_{e}$. $W_{e}$ is a learnable parameter, a part of our training end-to-end system. Since in our scope of experiments, each program $P$ is frequently broken down to fine-grained short repeatedly constituents $n_{i}$. Thus, normally a node $n_{i}$ often consists of a short sequence of tokens. For this reason, we employ a much simpler model, GRU [12], which is much simpler than state-of-the-art or frequently used recurrent models like Transformer [13], or LSTM [14]. GRU still employs the gate mechanism - a mechanism to model long dependency tokens interaction, similar to LSTM. But by dropping unnecessary forget gates as in LSTM, we reduce the number of parameters, improving training efficiency and less prone to overfitting in the case of our experiments.
Code tokens relation is not increasing order, left to right manner as in natural language but rather in both directions of appearance. We calculate the node embedding $\overline{x_{i}}$ of $n_{i}$ via each token embeddings $x_{i_{1}},x_{i_{2}},\cdots,x_{i_{L}}$ as follows:
$h_{i_{t}}=GRU(x_{i_{t}},h_{i_{t-1}})\quad\text{for}\,\,t=\{2,\cdots L\}\$ (1)
$\overline{x_{i}}=Average(h_{i_{1}},\cdots,h_{i_{L}})$ (2)
Element-wise average pooling operation $Average$ aggregates the state of a token to form final embedding $\overline{x_{i}}$ of the node $n_{i}$.
VI Dynamic Dependencies Learning
Figure 5: Dynamic Dependencies Learning
A successful execution in CFG is represented by a path $P$ with order starting from the “BEGIN” node denoted as $n_{0}$, and sequentially adds node $n_{i}$ until “EXIT” node. To check the coverability of one node $n_{j}$, we need to check if $n_{j}$ appears in $P$ or not. We can treat this problem as a binary classification task with classification label $t_{i}$=$1$ meaning node $n_{i}$ is covered and $t_{i}=0$ as node $n_{i}$ is not appearing in the execution path: $t_{i}\sim P(t|\{n_{j}\}_{j<i})$. However, not all the previous nodes in BFS are needed for the conditioning. But rather depends on the CFG itself. We represent the state of execution of node $n_{i}$ by $h_{i}\in\mathbb{R}^{n}$ with $h_{0}=\mathbb{0}$ at the beginning of the execution. Considering an execution path $P=\{n_{i_{1}},n_{i_{2}},\cdots,n_{i_{k}}\}$ of length $k$ with $i_{1}<i_{2}<\cdots<i_{k}$. The probability of appearance of a current node $i_{j}$ is conditioned on previous nodes: $t_{i_{j}}\sim P(t|n_{i_{1}},n_{i_{2}},\cdots,n_{i_{j-1}})$. One might consider using recent recurrent networks like traditional $RNN$ or $LSTM$ to model this conditioned distribution. With node embeddings $\overline{x_{1}},\overline{x_{2}},\cdots$ are computed in previous steps, we update the state $h_{i}$ by following equations:
$h^{*}_{i}$ $=f(h_{0},h_{1},\cdots,h_{i-1})$ (3)
$c^{*}_{i}$ $=f(c_{0},c_{1},\cdots,c_{i-1})$ (4)
$\text{in}_{i}$ $=\sigma(W_{\text{in}}\overline{x_{i}}+U_{\text{in}}h^{*}_{i}+b_{\text{in}})$ (5)
$\text{fg}_{i}$ $=\sigma(W_{\text{fg}}\overline{x_{i}}+U_{\text{fg}}h^{*}_{i}+b_{\text{fg}})$ (6)
$\text{op}_{i}$ $=\sigma(W_{\text{op}}\overline{x_{i}}+U_{\text{op}}h^{*}_{i}+b_{\text{op}})$ (7)
$\text{u}_{i}$ $=\text{tanh}(W_{\text{u}}\overline{x_{i}}+U_{\text{u}}h^{*}_{i}+b_{\text{u}})$ (8)
For computing the next state and memory cell $h_{i}$ and $c_{i}$:
$c_{i}$ $=\text{in}_{i}\odot u_{i}+\text{fg}_{i}\odot c^{*}_{i}$ (9)
$h_{i}$ $=\text{op}_{i}\odot\text{tanh}(c_{i})$ (10)
With function $f$ in (3) is aggregation function from previous states to summarize history information. For usual language modeling task, the function $f$ normally takes form of $f(h_{0},h_{1},\cdots,h_{i-1})=h_{i-1}$.
But considering the following characteristics from our problem, we propose an adaptation to the original LSTM which was originally used for language modeling $P(t_{i}|t_{0},t_{1},\cdots,t_{i-1})$
1. CFG edges aggeregation: in our CFG, a node $n_{i}$ has adjacency matrix considered only forward edges, denoted as $A_{\text{forward}}$. The aggregation function $f$ to be an average of adjacent nodes’ states, $f(h_{i})=A_{\text{forward}}H$ with $H=[h_{0},\cdots,h_{i-1},\mathbb{0},\cdots,\mathbb{0}]$
2. Forward and backward passing: in our CFG, a loop is broken down into condition node, body, and step node (e.g: $\text{counter}+=1$). The forward edges are just ordinary increasing order, $n_{i}\to n_{j}$ but what is special is an additional backward edge from the loop step node to the condition node. To propagate the information from the step node $n_{j}$ to condition node $n_{i}$ with $i<j$ but not to mix up the recurrent relations of execution. We update the state $h_{i}$ by $h_{j}$ for this special backward edge only by information of node embedding $\overline{x_{j}}$. By updating only the information but not the state at node $n_{j}$ which is not yet computed by forward order, we can combine both forward and backward propagation by only updating $h_{i}=LSTM(h_{j})$ with $i<j$ instead of bidirectional like $BiLSTM$.
3. Binary soft decision branching: we have processed a condition (including loop condition) node $n_{i}$ will only connect to two nodes $n_{j}$ and $n_{k}$ with $i<j,k$ by two forward edges. And each of these edges is the only incoming edge to $n_{j}$ and $n_{k}$. The original computation of node $n_{j}$’s hidden state is $h_{j}^{*}=A_{forward}H=[h_{0},\cdots,h_{i-1},\mathbb{0},\cdots,\mathbb{0}]=h_{i}$, similarly, $h_{j}^{*}=h_{i}$. With that, we allow information to pass to both possible branches while they are complemented in real code execution. Thus in order to model the branching behavior in the condition node, we will charge $A$ dynamically based on the current hidden state $h_{i}$. $A[i,j]=0$ if $Average(h_{i})\geq 0$ and $A[i,k]=0$ if $Average(h_{i})<0$ with assumption $j<k$. This will force the weights to adaptively produce reasonable $h_{i}$ to make a correct branch decision. A similar approach is also taken by [15], but the major difference is that they add additional parameter complexity to learn this soft dynamic branching while we focus more on the efficiency by setting branching conditions depending only on the current hidden state.
VII Coverage Prediction and Error Localization
VII-A Coverage Prediction
In this step, we use the hidden states $h_{i}$ learned from Step 3 (Section VI) to predict the code coverage. Each hidden state $h_{i}$ is passed through a linear layer followed by a sigmoid activation function to compute the coverage score. The score is then compared against a threshold $\alpha$ to classify whether a node is covered. Specifically, the process is defined as follows:
With $h_{i}$in $\mathbb{R}^{n}$ be the final hidden state for node $n_{i}$ after dynamic dependencies learning, we compute the coverage score:
$s_{i}=\sigma(W_{c}h_{i}+b_{c})$ (11)
where $W_{c}\in\mathbb{R}^{1\times n}$ and $b_{c}\in\mathbb{R}$ are the weights and bias of the linear layer, and $\sigma$ denotes the sigmoid activation function.
The coverage classification for node $n_{i}$ is determined by comparing $s_{i}$ to the threshold $\alpha$:
$\hat{t}_{i}=\begin{cases}1&\text{if }s_{i}\geq\alpha\\ 0&\text{otherwise}\end{cases}$ (12)
where $\hat{t}_{i}$ is the predicted coverage label for node $n_{i}$. A value of $\hat{t}_{i}=1$ indicates that node $n_{i}$ is predicted covered, while $\hat{t}_{i}=0$ indicates that it is predicted to be covered, while $\hat{t}_{i}=0$ indicates that it is predicted to be not covered.
To train our model, we use the Binary Cross-Entropy (BCE) loss function, which is suitable for binary classification tasks. The BCE loss for a single node $n_{i}$ is given by:
$\mathcal{L}_{i}=-\left[t_{i}\log(s_{i})+(1-t_{i})\log(1-s_{i})\right]$ (13)
where $t_{i}$ is the true label (1 if the node is covered, 0 otherwise) and $s_{i}$ is the predicted coverage score.
The total loss $\mathcal{L}$ over all nodes is the average of all losses:
$\mathcal{L}=\frac{1}{N}\sum_{i=1}^{N}\mathcal{L}_{i}$ (14)
where $N$ is the total number of nodes in the training set.
VII-B Runtime Error Detection and Localization
The underlying idea for runtime error detection is that code without runtime errors will terminate normally, covering both the BEGIN and EXIT nodes in the CFG. In contrast, buggy code will crash during execution, resulting in the EXIT node not being reached. Therefore, we focus on the coverage of the EXIT node to identify the presence of runtime errors.
One critical issue with existing models in predicting code coverage is the lack of continuity in the CFG. Discontinuity leads to gaps in the predicted execution path, making it difficult to accurately localize errors. CodeFlow, addresses this issue by consistently predicting a concrete, continuous path from the BEGIN to the EXIT node. This continuity ensures that the predicted execution flow closely follows the actual control flow of the program. Finally, to detect and localize runtime errors, we analyze the predicted code coverage as follows:
  • Runtime Error Check: If CodeFlow predicts EXIT node as a covered node, the code is likely free of runtime errors. If it does not, we infer that the code has crashed.
  • Error Localization: In the buggy code, the furthest node reached without an outgoing edge is identified as the crash point, indicating the location of the runtime error.
By leveraging the continuity and comprehensive path prediction capabilities of CodeFlow, we improve the reliability of error detection and localization.
VIII Empirical Evaluation
For evaluation, we seek to answer the following questions:
RQ1. [Code Coverage Prediction Accuracy]: How well does CodeFlow predict code coverage for (in)complete code in Python?
RQ2. [Runtime Error Detection Accuracy]: How well does CodeFlow detect runtime errors in (in)complete code?
RQ3. [Runtime Error Localization Accuracy]: How accurately does CodeFlow locate the lines with a runtime error?
RQ4. [Usefulness in Fuzz Testing]: How useful does CodeFlow support fuzz testing in detecting runtime errors for (in)complete code snippets?
IX Code Coverage Prediction Accuracy (RQ1)
IX-A Data Collection, Baselines, Procedure, and Metrics
IX-A1 Datasets
For training, we utilize a comprehensive dataset specifically curated for code coverage prediction. Our primary dataset, CodeNetMut, is derived from Liu et al. [9]. This dataset was created by crawling and generating mutations based on submissions to competitive programming problems from the CodeNet dataset [16]. CodeNetMut contains nearly 20,000 Python files. After excluding those that failed execution by python-trace or CFG construction by python-graphs, we were left with 8,216 Python code snippets.
However, CodeNetMut lacks a sufficient number of Python files with extensive conditional statements, which are crucial for training the model on the behavior of conditional branching. To address that, we supplemented CodeNetMut with an additional dataset generated using Gemini-API. This synthetic dataset comprises approximately 11,668 Python code snippets, each featuring diverse and complex types of statements. For a snippet, we generate the CFG in which we track the nodes, modeling individual statements or blocks of code, forward edges (representing the normal control flow from a statement to another), and backward edges (representing the control flow from the last line of a loop back to the loop node.
To build the ground truth in training, we use the trace library from Python to record which lines of code are executed when running each snippet and record the final code coverage.
IX-A2 Baselines and Procedure
We compare CodeFlow with the following approaches:
1. CodeExecutor [9]: primarily predicts execution traces. It leverages the transformer-based UniXcoder model, which is trained via the data including source code, input values, and the full execution traces with values at each execution step.
2. CFGNN [17]: originally designed for detecting condition-related bugs via CFGs. We modified CFGNN by retaining its main architecture but altering the final linear layer to output a list of scores for each node, instead of a single node, allowing it to predict coverage across multiple nodes.
3. OpenAI GPT-4o (gpt-4o), Anthropic Claude (claude-3.5-sonnet), and Google Gemini (gemini-1.5-flash): We used several LLMs as baselines. We follow Tufano et al. [18] to design the prompt to GPT-4 to get the code coverage.
IX-A3 Evaluation Metrics
To evaluate the models, we use the following key metrics.
+ Exact Matching (EM): This metric counts the number of times when the entire predicted sequence of statements exactly matches the target sequence of true coverage, representing the model’s capability to predict the executed statements.
+ Branch Coverage Matching (BC): This metric counts the number of times a model correctly predicts the branch coverage (at a condition node in CFG), assessing a model’s prediction on conditions and loops.
+ Precision (P), Recall (R), F1-Score (F1): Precision is the ratio of correctly predicted executed nodes (true positives) to the total number of nodes predicted as executed (true positives + false positives). Recall is the ratio of correctly predicted executed nodes (true positives) to the total number of actually executed nodes (true positives + false negatives). The F1-Score is the harmonic mean of Precision and Recall.
IX-B Empirical Results
TABLE I: Code Coverage Prediction Comparison (RQ1).
Model EM (%) BC (%) P R F1
CodeEx 18.83 31.34 0.94 0.47 0.63
CFGNN 45.53 76.56 0.92 0.91 0.92
LLMs
   Gemini 56.17 74.96 0.87 0.97 0.92
   Claude 64.94 77.30 0.96 0.94 0.95
   GPT-4 68.13 78.75 0.96 0.96 0.96
CodeFlow 75.24 87.88 0.97 0.97 0.97
As seen in Table I, CodeFlow outperforms existing models across all key metrics.
IX-B1 Exact-Matching
CodeFlow achieves an exact matching accuracy of 75.24%, which is higher than the best-performing LLM, GPT-4o, at 68.13%. Notably, CodeFlow accomplishes this with far fewer parameters (1.3 million in total) compared to LLMs like GPT-4o (over 1 trillion) and Claude (175 billion), which suggests a more efficient architecture for this task and underscores its practicality for the scenarios where computational resources may be limited.
Figure 6: Code coverage prediction visualization for different models with red nodes indicate coverages. For CFGNN and CodeFlow, the nodes’ shades correspond to values from 0 to 1, representing the predicted coverage scores.
IX-B2 Branch-coverage Matching
CodeFlow achieves 87.88% correctness, outperforming the next best baseline, GPT-4o, by nearly 10%. This margin highlights the efficacy of using CFGs to capture the complex behaviors of loops and conditional branches. With CFG modeling the intricate decision points within code, CodeFlow understands the cumulative effect of variable changes over multiple iterations, leading to more accurate branch predictions. In contrast, models like CodeExecutor fail in this aspect because they do not adequately handle dynamic execution changes, leading to predictions that do not align well with actual execution paths involving loops and conditions.
CodeFlow also excels in precision, recall, and F1-score metrics, achieving 97% in all three categories. These values are slightly higher than those of the best LLMs (Claude and GPT-4o), which both scored 96% in precision and recall. The high precision indicates that CodeFlow is very effective in identifying executed lines without falsely predicting non-executed lines as executed (minimizing false positives). The recall score reflects the model’s ability to capture all executed lines without missing any (minimizing false negatives). Regarding the LLMs, note that half of the test dataset comprises solutions from the CodeNet Project published online since 2022, potentially giving them an advantage if they were trained on this data. Despite this, CodeFlow still outperforms them.
IX-B3 Continuity in Predicting Code Coverage
Continuity means that a model should predict nodes or lines of code that are close together or logically lead to each other, rather than randomly jumping to distant nodes while ignoring intermediate nodes in between. This is a common issue observed with existing methods. CodeExecutor relies heavily on exact execution traces during training, which causes it to miss the dependencies between lines of code. As seen in Fig. 6a, CodeExecutor skips one node and jumps directly to the next node, resulting in a discontinuous path in the CFG.
Similarly, CFGNN transmits information equally through all possible paths in the CFG, rather than focusing on the correct execution paths. This approach often leads to a misunderstanding of the continuity in code coverage tasks. Fig. 6b clearly shows the heatmap of predicted scores for each node, highlighting CFGNN’s shortcomings in coverage prediction.
LLMs, such as GPT-4o, also face challenges in maintaining continuity. They sometimes fail to understand the dynamic relations between lines and may skip essential steps in sequences, e.g., the elif in an if-elif-else structure. This is depicted in Fig. 6c, where the model predicts nodes without capturing the logical sequence. This occurs because LLMs rely primarily on next-token probability, predicting the most likely subsequent token based on the preceding context. This approach can result in misunderstandings of the code structure, especially concerning dynamic behaviors and state changes across iterations. Thus, while having high precision, recall, and F1-score, they have lower exact-matching accuracies.
CodeFlow, while similar to CFGNN in using CFGs, is designed to address this issue. By emphasizing the correct path through CFGs and ensuring that information is passed predominantly along the actual execution paths, CodeFlow maintains continuity in its predictions (Fig. 6d). The model predicts sequences of executed lines that are connected, following a coherent path from the beginning to the end.
IX-B4 Incomplete Code
To our knowledge, there is no existing runtime error dataset of incomplete Python code snippets with the respective complete versions (to build the ground truth). Thus, to test its capability with incomplete code, specifically when built-in and external libraries are not imported, we trained and tested CodeFlow on a dataset where we removed all import statements of built-in/external libraries or method/class declarations. This ensures the code’s semantic integrity remains intact including keeping all function/API calls within the code. Despite the absence of explicit import statements, CodeFlow achieved similar results as in Table I (not shown) because it learned the semantic meaning of these libraries during training, allowing it to understand and predict code behavior based on the nodes that use the library functions. Similar performance was observed in LLMs (GPT-4o, Claude, and Gemini), as their extensive pre-training includes library semantics. CodeExecutor, which tracks intermediate values in the execution trace, and CFGNN, which models control flow, also performed well without import information.
X Runtime Error Detection Accuracy (RQ2)
In this study, we assess the ability of CodeFlow to predict whether a given code snippet contains a runtime error.
X-A Data Collection and Evaluation Metrics
For the purpose of evaluating runtime error detection accuracy, we utilized a dataset in addition to the one in RQ1. Specifically, we employed the FixEval dataset, which comprises 2,066 unique problems with a total of 277,262 submissions of Python code snippets. From this dataset, test cases were obtained for 800 problems from the CodeNet dataset [16]. Each of these Python snippets, when executed with its respective input, leads to a runtime error.
After filtering, we obtained 6,437 submissions across the 800 problems, which we added to our dataset. This combined dataset, referred to as the Complete Runtime-Error Dataset, includes both code snippets that terminate normally and those that encounter runtime errors.
To evaluate CodeFlow in detecting runtime errors, in addition to the metrics in RQ1, we use Runtime Error Detection Accuracy (EDA) to measure the accuracy of a model correctly predicting if a code snippet has an error.
X-B Empirical Results
TABLE II: Runtime Error Detection Comparison (RQ2).
Model EDA (%) P R F1
CFGNN 76.71 0.51 0.89 0.65
LLMs
   Claude 77.98 0.98 0.89 0.93
   GPT-4o 69.24 0.71 0.99 0.83
CodeFlow 97.51 0.96 0.94 0.95
As presented in Table II, Claude achieved the highest precision score of 0.98, indicating its strong ability to correctly identify runtime errors when they are present. High precision means Claude makes very few false positive predictions, thus demonstrating its reliability in accurately pinpointing real runtime errors. However, Claude’s recall score of 0.89, while still respectable, is lower than that of GPT-4o. This suggests that Claude may miss some actual runtime errors, indicating that it is more conservative in its error detection.
In contrast, GPT-4o achieved the highest recall score of 0.99, which highlights its effectiveness in identifying nearly all actual runtime errors. This high recall shows GPT-4o’s thoroughness in detecting errors. However, GPT-4o’s precision score of 0.71 indicates a higher rate of false positives compared to Claude, meaning it sometimes incorrectly flags non-buggy code snippets as erroneous. This suggests that GPT-4o could have more false alarms.
CFGNN shows a more moderate performance with a runtime error detection accuracy of 76.71%, a precision of 0.51, a recall of 0.89, and an F1-score of 0.65. This implies that while CFGNN can detect errors, it struggles to accurately discriminate between erroneous and non-erroneous code snippets, leading to many false alarms. This performance is likely due to CFGNN’s approach of transmitting information equally through all possible paths in the CFG, which might result in overestimating the likelihood of errors.
As seen, CodeFlow exhibits a balanced performance with high scores across all metrics: a precision of 0.96, a recall of 0.94, and an F1-score of 0.95. This balance indicates that CodeFlow not only accurately detects a high proportion of actual runtime errors but also minimizes false positives. The overall detection accuracy of 97.51% underscores CodeFlow’s superior capability in statically identifying runtime errors without execution.
Notably, the performance of all models remained relatively stable even when tested on incomplete code snippets (not shown). This indicates that the models, including CodeFlow, can understand the semantic meaning of the removed library, generalize well, and maintain high detection accuracy.
We did not use CodeExecutor as a baseline for runtime error detection because it is only trained on datasets with full execution traces that lack instances of crashes and runtime errors. Consequently, it always provides the execution trace and intermediate values until the end of execution, rather than detecting or stopping at crash points, failing to detect errors.
XI Runtime Error Localization Accuracy (RQ3)
After detecting whether a snippet contains errors, the next step is to localize the specific lines that raise these errors.
XI-A Data Processing and Evaluation Metrics
We evaluated all models on approximately 1,300 different buggy code snippets from the FixEval dataset [19]. In addition to the Complete Runtime Error Dataset in RQ2, we created the Incomplete Runtime Error Dataset by removing all import statements and external file references from each snippet.
In this section, we focus on a new metric, Error Localization Accuracy, which measures the number of times the predicted buggy line matches the actual buggy line. Based on the results from Section X, we observed that Claude performed the best in detecting runtime errors. Therefore, for this experiment, we use Claude as the main baseline to compare with our model.
XI-B Empirical Results
TABLE III: Error Localization Accuracy Comparison (RQ3)
Metric Claude CodeFlow
Complete Incomplete Complete Incomplete
Err Loc Acc 60.20% 59.41% 72.22% 70.37%
The results in Table III show that CodeFlow significantly outperforms Claude in runtime error localization accuracy. For the Complete Runtime Error Dataset, CodeFlow achieved an accuracy of 72.22%, compared to Claude’s 60.20%. This demonstrates CodeFlow’s superior ability to accurately pinpoint the exact lines causing runtime errors for complete code. Similarly, for the Incomplete Runtime Error Dataset, CodeFlow maintained a high bug localization accuracy of 70.37%, while Claude’s performance slightly dropped to 59.41%. This consistency highlights CodeFlow’s effectiveness even when external library imports are removed, indicating that the model can still understand and trace the flow in the code accurately.
XI-B1 HeatMap Visualization
In addition to Table III, further analysis of the heatmap visualization in Fig. 7 for an example provides insights into CodeFlow’s effectiveness in handling buggy code. In the heatmap, we observe that the scores of nodes do not significantly drop in buggy code, unlike in non-buggy code where the score for nodes not in the covered path drops very low (indicated by white color in Fig. 6d). In buggy code, the crash point reduces the score, but the nodes in the likely-correct path to the EXIT node still retain high scores.
Figure 7: Heatmap visualization of node scores in buggy code.
To further enhance error localization accuracy, we experimented with increasing the $\alpha$ value in Section VII-A to classify node. By filtering out more non-covered nodes, we observed improved accuracy as shown in Table IV.
TABLE IV: Error Localization Accuracy with Different Alpha Values (RQ3)
Alpha Value Error Localization Acc
$\alpha=0.5$ 72.22%
$\alpha=0.7$ 74.82%
$\alpha=0.9$ 77.40%
$\alpha=0.95$ 78.31%
As we increase the alpha value from 0.5 to 0.95, the bug localization accuracy improves, reaching 78.31% at $\alpha=0.95$. This indicates that by setting a higher threshold, CodeFlow becomes more effective at filtering out non-relevant nodes, thereby enhancing its ability to identify the buggy lines.
Additionally, to further demonstrate the effectiveness of CodeFlow in runtime error detection, we have compiled a list of the top 10 runtime errors that our model can successfully identify. Table V highlights these common runtime errors along with their corresponding error messages.
TABLE V: Top 10 Runtime Errors Detected by CodeFlow
Runtime Error Error Message
Operand Type Mismatch unsupported operand type(s) for ** or pow(): ’str’ and ’int’
Comparison Error ’¡’ not supported between instances of ’list’ and ’int’
Object Not Callable ’int’ object is not callable
Non Iterable Type ’int’ object is not iterable
Invalid Argument Type list indices must be integers or slices, not str
TypeError ’float’ object cannot be interpreted as an integer
Type Specific Operation can’t multiply sequence by non-int of type ’str’
Non Subscriptable ’int’ object is not subscriptable
Attribute Error object of type ’int’ has no len()
NoneType Subscripting ’NoneType’ object is not subscriptable
XII Usefulness in Fuzz Testing (RQ4)
For this study, we evaluated the usefulness of CodeFlow in supporting fuzz testing to detect and localize runtime errors in incomplete/non-executable code.
XII-A Fuzz Testing Procedure
The fuzz testing procedure consists of three main steps:
  • Input Generation: We employed the LLM Claude to generate inputs that are likely to raise runtime errors in the provided code snippets.
  • Runtime Error Detection: The code snippet with the generated inputs encoded in the assignment statements at the beginning was fed into CodeFlow. The model processed the code to determine whether it contained a runtime error and, if so, localized the specific buggy statement.
  • Feedback loop: If no runtime error is detected in Step 2, the process enters a feedback loop. The inputs from Step 1 that failed to raise an error were fed back into the LLM to regenerate new inputs. This process continues until a runtime error is discovered or the time limit is exceeded.
XII-B Empirical Results
TABLE VI: Runtime Error Detection Comparison (RQ4)
Metric Claude Fuzz Testing w/ CodeFlow
30s 60s 120s
#runtime error detected 32/50 44/50 46/50 47/50
Error Localization Acc 42.27% 49.53% 50.00% 42.61%
To evaluate the effectiveness of CodeFlow in supporting fuzz testing, we conducted experiments on 50 buggy Python code snippets from the FixEval dataset in which we removed all input variables and import statements. The results, summarized in Table VI, show that while Claude by itself detected 32 runtime errors, incorporating its fuzzing with CodeFlow detected 44 errors within 30 seconds (37.5% relative improvement), 46 errors within 60 seconds, and 47 errors within 120 seconds.
Additionally, the Error Localization Accuracy with CodeFlow was consistently higher than that of Claude, as demonstrated in RQ3. It is particularly challenging for LLMs like Claude to detect the correct line containing bugs in incomplete code due to the lack of inputs, making it difficult for Claude to reason about code execution and runtime behaviors.
The significant enhancement from CodeFlow is particularly valuable in incomplete code, where direct execution is infeasible due to the nature of incomplete code. By using Claude to generate likely inputs and CodeFlow to predict runtime errors without requiring external library setups, we effectively address this challenge. This integrated approach not only increases the detection rate of runtime errors but also provides reasonably accurate localization of the specific lines causing these errors. Other advanced LLM-based fuzzers, such as Fuzz4All [20], could be used with our framework.
XIII Related Work
Predictive Execution
CodeExecutor [9] was pre-trained on a dataset including source code, input values, and execution traces with values at each step. Its transformer learns to convert input and source code into execution traces. Ding et al. [10] introduce TRACED, an execution-aware pre-training strategy using a mix of source code, executable inputs, and execution traces. We did not compare with TRACED since it works only for C. LExecutor [21] predicts and injects missing values to execute arbitrary (in)complete code. It still requires execution. TraceFixer [22] is trained using buggy code, execution traces, desired values, and expected bug-fixed code. Bieber et al. [15] learn to execute on CFGs. They introduce additional parameter complexity to learn soft dynamic branching, whereas we set branching conditions based only on the current hidden state. Tufano et al. [18] prompt to LLM to return the code coverage.
ML-based Fault Localization
Early neural network-based fault localization (FL) approaches [23, 24, 25, 26, 27] primarily rely on test coverage data and struggle to distinguish between elements executed by failed tests and actual faulty elements [27]. In contrast, recent deep learning-based approaches such as GRACE [28], DeepFL [29], CNNFL [30], and DeepRL4FL [31] have demonstrated improved performance. GRACE introduces a novel graph representation for methods and learns to rank faulty methods.
Earlier learning-based FL techniques include MULTRIC [32], TrapT [27], and Fluccs [33]. Automated program repair approaches [34, 35] focus on locating and fixing bugs. The Hercules APR tool [35] can identify multiple buggy hunks. FixLocator [36] detects co-fixing locations, and TRANSFER [37] utilizes deep semantic features and transferred knowledge from open-source data to enhance FL. CodeT5-DLR [38] introduces an end-to-end pipeline on LLMs to detect, localize and repair bugs in sequential order.
XIV Conclusion & Future Work
Current code models often overlook dynamic dependencies between lines of code, focusing only on plain text or correct execution traces. To address this, we introduce CodeFlow, an approach that predicts code coverage and detects runtime errors by learning both static and dynamic dependencies. CodeFlow uses CFGs and a Gated recurrent unit (GRU) network to represent execution paths and learn vector embeddings for CFG nodes. It also leverages execution traces via CFG to capture dynamic dependencies among statements. Our evaluation shows that CodeFlow significantly improves coverage prediction accuracy and runtime error localization, outperforming state-of-the-art models.
While CodeFlow has shown promising results, there are several avenues for future research and improvements:
  • Extending to other programming languages: Currently, CodeFlow has been tested primarily on Python code. Future work could involve adapting the model to support other popular programming languages, expanding its applicability.
  • Integration with development environments: Exploring ways to integrate CodeFlow into integrated development environments (IDEs) could provide real-time code coverage and error detection capabilities to developers.
  • Handling larger codebases: Investigating techniques to handle larger, more complex codebases and entire software projects would enhance its practical utility.
  • Combining with LLMs: Exploring hybrid approaches that leverage the strengths of both CodeFlow and LLMs could potentially lead to even more powerful code analysis tools.
  • Dynamic code analysis: Extending CodeFlow to perform more complex dynamic analysis tasks, such as predicting performance bottlenecks or potential vulnerabilities.
  • Continuous learning: Developing mechanisms for CodeFlow to continuously learn and improve its predictions based on feedback could enhance its effectiveness.
In conclusion, CodeFlow represents a significant step forward in predicting runtime program behavior without execution. By effectively modeling both static and dynamic code dependencies, it opens up new possibilities for code analysis and software quality assurance. As we continue to refine and expand this approach, we anticipate CodeFlow and similar graph-learning-based models will play an increasingly important role in the future of the static prediction of runtime behaviors. Our data and code is available in our website [39].